### v1

# GimbalDiffusion: Gravity-Aware Camera Control for Video Generation

Abstract
Recent progress in text-to-video generation has achieved remarkable realism, yet fine-grained control over camera motion and orientation remains elusive. Existing approaches typically encode camera trajectories through relative or ambiguous representations, limiting explicit geometric control. We introduce GimbalDiffusion, a framework that enables camera control grounded in physical-world coordinates, using gravity as a global reference. Instead of describing motion relative to previous frames, our method defines camera trajectories in an absolute coordinate system, allowing precise and interpretable control over camera parameters without requiring an initial reference frame. We leverage panoramic $360^{\circ}$ videos to construct a wide variety of camera trajectories, well beyond the predominantly straight, forward-facing trajectories seen in conventional video data. To further enhance camera guidance, we introduce null-pitch conditioning, an annotation strategy that reduces the model’s reliance on text content when conflicting with camera specifications (e.g., generating grass while the camera points towards the sky). Finally, we establish a benchmark for camera-aware video generation by rebalancing SpatialVID-HQ for comprehensive evaluation under wide camera pitch variation. Together, these contributions advance the controllability and robustness of text-to-video models, enabling precise, gravity-aligned camera manipulation within generative frameworks.
“A serene aerial view captures a boat drifting down a dark river, framed by lush greenery and shimmering water under a peaceful sky.”
Refer to caption Refer to caption Refer to caption
“A sleek, modern home with high ceilings, natural light, and contemporary furnishings features a man in a blue suit in its open-plan living area.”
Refer to caption Refer to caption Refer to caption
Reference Ours AC3D [bahmani2025ac3d]
Figure 1: Reproducing real scenes using our representation. From reference frames from a driving video (left), GimbalDiffusion can be controlled to replicate its camera viewpoint with respect to gravity (center). In contrast, existing methods (right) do not accurately match the absolute reference pose.
1 Introduction
Text-to-video generation is rapidly advancing, offering astonishing photorealism and new creative workflows. Yet, the ability to control generated video content—especially camera movement and trajectory—remains limited. Early approaches rely solely on textual input which does not provide fine-grained control, such as specifying geometric parameters for camera paths and angles.
Recent methods [he2024cameractrl, he2025cameractrl, bahmani2025vd3d, bahmani2025ac3d, bai2025recammaster, wan2025] address this limitation by conditioning video models with explicit camera trajectories, frequently encoded as Plücker rays. However, these representations are inherently ambiguous and lack the precision of real camera systems. Typically, camera extrinsics used for Plücker conditioning are defined relative to the first frame, making it impossible to explicitly control camera orientation with an absolute reference within the environment (e.g., with respect to gravity).
In this work, we introduce GimbalDiffusion, a novel gravity-centric approach to data annotation that overcomes this challenge and enables precise, absolute user control over generated camera trajectories, as illustrated in LABEL:fig:Teaser. The name reflects the role of a gimbal—a device that stabilizes orientation in physical space—mirroring our use of gravity as a global reference to stabilize and disambiguate camera rotation. To resolve rotational ambiguity, we leverage geometrically calibrated 360° panoramic videos. By sampling camera crops over the sphere, we can remove the camera bias in human-captured datasets, allowing rare camera shots such as Dutch angles (45° roll) or extreme pitches to be sampled. This alignment enables the generation of cinematic shots, such as low- and high-angle perspectives, which are uncommon in regular video datasets. The effectiveness of this approach is demonstrated in fig. 1, where our method produces frames with camera views that more closely match the intended trajectories than previous methods.
To design our dataset from panoramic $360^{\circ}$ videos, original translation paths are preserved while camera rotations are randomly sampled. This lets our model learn a wide variety of camera trajectories, far beyond the predominantly straight, forward-facing paths seen in conventional video data.
We further observe that text and camera angles are inherently entangled, and that care must be taken to ensure that appropriate control is achieved. For example, when asked to generate a video of a “green field under a blue sky” with a camera looking down, a model must learn to render grass, even if not specified in the prompt. To this end, we propose a null-pitch conditioning approach, where captioning is performed on forward-facing crops rather than the actual camera point of view.
Finally, we propose a new evaluation benchmark for video generation with extreme camera angles. To do so, we use a large-scale dataset of high-quality internet videos, SpatialVID-HQ [wang2025spatialvidlargescalevideodataset], and rebalance the camera pitch to obtain diverse test samples. We will release this test dataset upon publication.
In summary, we make the following contributions:
  • A pipeline for generating diverse camera trajectories from panoramic 360° video, providing exhaustive coverage of possible camera orientations;
  • A strategy named null-pitch conditioning for disentangling text from absolute camera angle;
  • A novel benchmark for video generation with extreme camera angles.
Figure 2: Training data pipeline. (a) We extract the camera poses from a dataset of 360° panoramic videos in equirectangular format (see Sec. 3.2). (b) We randomly sample a camera rotation trajectory on the sphere, from which we generate field of view masks in equirectangular format for both the projected image and the null-pitch conditioning (see Sec. 3.3). (c) We integrate the two previous steps by combining the sampled rotation with the estimated camera poses. This process creates a dataset of perspective images with their camera poses, along with null-pitch reference images. The main advantage of this approach is the ability to sample camera viewing directions across the entire sphere, in contrast to using natural human-captured videos, which are heavily biased towards null roll and pitch. Additionally, we generate a text description for both the perspective video and the null-pitch video using a VLM.
2 Related Work
Video generation.
Generative video modeling has evolved rapidly, from early GAN-based methods that produced short, low-resolution clips [tulyakov2018mocogan, li2018video, Acharya2018] to large-scale diffusion-based architectures capable of generating temporally coherent, semantically consistent videos [Weng2024]. Recent open-source models such as CogVideoX [yang2024cogvideox] and WAN [wan2025] push the frontier of video length, resolution, and controllability while proposing both text-to-video and image-to-video variants.
Editing generated videos remains challenging: precise modification to content, viewpoint, or lighting is difficult once frames are synthesized. To address this, several control strategies have emerged, introducing user-guided conditioning based on edge maps, depth, sketches, or explicit 3D geometry for structural control [chen2023controlavideo, Yenphraphai2024ImageSculpting, chang2024magicpose, hu2023animateanyone], and illumination-aware guidance for lighting control [kocsis2024lightit, Magar_lightlab]. These mechanisms allow precise control over motion, composition, and scene appearance, substantially improving the usability and compositional fidelity of generative video systems.
Camera control methods.
Beyond content and appearance, viewpoint control has become key in generative video. A common representation for such control is Plücker rays, described in Sec. 3.1. This formulation enables differentiable, geometrically consistent camera manipulation.
In static image generation, PreciseCam [bernal2025precisecam] introduces absolute control over camera pose by conditioning on roll, pitch, field of view, and lens distortion alongside textual prompts. For image-to-video generation, methods such as MotionCanvas [xing2025motioncanvas], Diffusion-as-Shader [gu2025das], and TrajectoryCrafter [mark2025trajectorycrafter] employ point cloud conditioning to maintain spatial coherence while navigating a scene from an initial reference image. While effective, these approaches depend on the depth and pose estimation quality of the initial reference frame, which can restrict flexibility and generalization.
In text-to-video synthesis, AC3D [bahmani2025ac3d] allows user-defined camera trajectories but conditions motion relative to the first frame only, leaving global rotational and translational ambiguities unresolved. Similarly, Li et al. [li2025cameras] manipulate positional encodings to control viewpoint changes, yet still lack a consistent absolute reference such as gravity.
These works underscore the need for representations that enable precise, absolute camera control and scalable trajectory generation across space and time. Our approach addresses this gap by aligning camera orientation with gravity and anchoring motion within a consistent world frame, advancing controllable text-to-video synthesis.
Camera pose estimation.
Estimating accurate camera poses from images is a long-standing problem in 3D vision. Classical structure-from-motion pipelines such as VisualSFM [wu2011visualsfm], OpenMVG [moulon2016openmvg], and COLMAP [Schonberger2016COLMAP] reconstruct camera intrinsics and extrinsics through feature matching, triangulation, and global bundle adjustment. More recent learning-based approaches jointly infer scene geometry and pose from data: VGGT [wang2025vggt] leverages transformer-based global geometry tokens for pose regression, while DUSt3R [wang2024dust3r] and MASt3R [leroy2024grounding] predict dense correspondence fields to recover metrically scaled 3D structure. MegaSAM [Jiang2024MegaSAM] extends these ideas to foundation scale, achieving state-of-the-art alignment accuracy.
Despite their accuracy, these systems typically operate in arbitrary coordinate frames and lack physical grounding in gravity or a global world reference, limiting their absolute interpretability. To mitigate this, several works estimate the gravity direction from single images [jin2022PerspectiveFields, veicht2024geocalib, xian2019uprightnet]. However, acquiring datasets with gravity annotations often requires the use of auxiliary sensors, such as accelerometers, or additional semantic information about the scene. Examples include EDINA [Do_2022_EgoSceneMSR], Horizon Lines in the Wild [workman2016hlw], and the KITTI Horizon dataset [kluger2020temporally].
To circumvent such constraints, we leverage geometrically calibrated 360° panoramic videos from WEB360 [huang2025vipe] as backbone for our data generation. By combining the up vector derived from the panoramic video with SfM-estimated intrinsics and extrinsics, we obtain a consistent absolute reference for gravity-aware camera control.
Figure 3: Training data samples from our data augmentation pipeline, capturing a highly diverse set of rotation trajectories from 360° videos. Both the trajectory and the prompt are generated on-the-fly in our dataloader. Each color in the graph corresponds to a video above.
3 Method
We aim to train a camera-conditioned model on precise, gravity-referenced camera viewpoints. We refer to this global gravity reference as an absolute reference, in contrast to existing relative methods that describe camera motion with respect to a previous frame. Because our absolute reference is aligned with gravity, pitch and roll are well defined, whereas yaw remains unconstrained. Throughout the paper, absolute rotation therefore denotes pitch and roll expressed in this shared gravity-aligned reference.
However, no large-scale video dataset provides the gravity annotations we need to train such a model. This section outlines our pipeline for obtaining such data, summarized in fig. 2. In fig. 2 (a) we extract camera pose annotations from a dataset of 360° panoramic videos, as detailed in Sec. 3.2. Next, in (b), we sample a camera trajectory as described in Sec. 3.2. Finally, in (c), we combine the two previous steps to produce our dataset. We repeat the same procedure for an additional version of each trajectory with null pitch and roll, described in Sec. 3.3. Before detailing the full pipeline, the following section explains the standard representation to encode camera poses in neural-network-compatible buffers.
3.1 Representing cameras
Given a camera extrinsic matrix $\mathbf{E}_{f}\!=\![\mathbf{R}_{f}|\mathbf{t}_{f}]$ and intrinsic matrix $\mathbf{K}_{f}$, a common strategy for conditioning video generation models on viewpoint is to use Plücker rays [kant2024spad, he2024cameractrl, he2025cameractrl, bahmani2025vd3d, bahmani2025ac3d]. For a conditioning frame $f$, we can define a per-pixel map of Plücker rays as $\mathbf{p}_{f,u,v}\!=\!(\mathbf{t}_{f}\times\mathbf{d}_{f,u,v}^{\prime},\mathbf{d}_{f,u,v}^{\prime})$, where $\mathbf{t}_{f}$ denotes the camera origin in world coordinates, and $\mathbf{d}_{f,u,v}$ is the direction vector from the camera center toward the center of pixel $(u,v)$. Each pixel is thus represented as a 6D vector describing a ray in 3D space. The ray direction can be computed as $\mathbf{d}_{f,u,v}\!=\!\mathbf{R}_{f}\mathbf{K}_{f}^{-1}[u,v,1]^{T}+\mathbf{t}_{f}$, with $\mathbf{d}_{f,u,v}^{\prime}$ denoting the normalized direction vector $\mathbf{d}_{f,u,v}$.
3.2 Creating varied camera paths from 360° videos
Most video datasets exhibit a significant bias towards forward-looking views, where the horizon remains level or close to the center of the image, as shown in fig. 4. However, such “normal” viewpoints restrict the expressive range that a filmmaker might wish to achieve in a storyboard. For instance, these datasets often lack cinematic shots, such as top-down, low-angle, or Dutch-angle shots. As a result, generative video models trained on such data struggle to produce these types of cinematic shots, since they fall outside the training distribution.
To address this limitation, we propose generating synthetic camera trajectories by sampling from 360° videos and rectifying them so that the horizon position is known throughout each sequence.
Camera poses from SfM.
From a 360° video, we compute the relative motion between each frame by extracting six perspective crops (front, back, left, right, top, bottom), and processing the resulting video with structure-from-motion (SfM). This results in a set of camera poses $\mathbf{E}_{\text{SfM},f}$ at each frame $f$. Taking the first frame of the sequence $f=0$ as reference, we can then compute the relative motion
$\mathbf{E}_{\text{rel},f}$ $=\mathbf{E}_{\text{SfM},0}^{-1}\mathbf{E}_{\text{SfM},f}\;.$ (1)
Sampling random camera paths.
We can generate several perspective videos from a 360° sequence by sampling random camera rotation paths. We outline our approach in Algorithm 1. We first sample random pitch, roll, and yaw angles uniformly to obtain an initial viewpoint. Next, 1 to 4 additional random rotations are sampled and distributed across the sequence length. Those rotations are applied sequentially to produce smooth and varied motion paths.
We further introduce variations in the camera field of view (FoV). For each sequence, we sample either 1, 2, or 3 keyframes at random timestamps, with an FoV ranging from 35° to 100°. The FoV values are then interpolated using cubic splines with randomized derivative boundary conditions. Examples from our sampling are shown in fig. 3. This approach provides broad coverage of the possible viewpoints, as shown in fig. 4. Notably, this sampling slightly biases viewpoints towards the poles, which we found to better support model stability and representation of more extreme viewpoints.
Gravity annotation.
To anchor the videos within a shared global reference frame, we align their rotation with gravity. We are able to compute the absolute, gravity-aligned camera pose for each frame as
$\mathbf{E}_{\text{abs},f}=\varphi(\mathbf{R}_{\text{pano},0})\mathbf{R}_{\text{pano},f}\mathbf{E}_{\text{rel},f}\;,$ (2)
where $\varphi(\cdot)$ is a function that preserves only the roll and pitch angles (removes the yaw) of the first sampled camera pose. After this adjustment, the transformed pose $\mathbf{E}_{\text{abs},f}$ has no rotation about the global up-vector axis. This function $\varphi(\cdot)$ is obtained by computing the gravity direction in camera coordinates and inverting it to recover the corresponding up-vector, see the supplementary for the complete definition. This process establishes a gravity-aligned coordinate system that is consistent across all videos, allowing rotations to be defined within a common global frame. Consequently, each sequence is initialized at null yaw, reducing ambiguity and easing training on this representation.
Figure 4: Comparison of parameter distribution between our sampling and a typical video dataset for camera control (RealEstate10K [RealEstate10K]). Our sampling produces videos with more rotational motion, as measured by total angular distance (the sum of absolute angular differences between consecutive camera poses), and greater diversity in Euler angles (pitch, roll, yaw). All values on the $x$-axis are expressed in degrees.
Up (+90°) Forward (0°) Down (-90°)
PreciseCam
Refer to caption Refer to caption Refer to caption
w/o null pitch
Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption
Prompt: A prairie with grass and flowers, and a beautiful blue sky with few clouds.
Figure 5: Entanglement between prompt and camera pitch. Without a careful captioning strategy, diffusion models may ignore the camera conditioning when the prompt semantics conflict with it. Our null-pitch conditioning mitigates this issue and preserves accurate camera control.
Algorithm 1 Random camera path sampling
[…]
6:for $i=1$ to NN do
7:  $d_{t}\sim\text{Uniform}(0,1)$ $\triangleright$ Random rotation duration
8:  $t_{s}\sim\text{Uniform}(0,1-d_{t})$ $\triangleright$ Random start time
9:  $t_{e}\leftarrow t_{s}+d_{t}$ $\triangleright$ End time
10:  $\mathbf{a}\sim\mathbb{S}^{2}$ $\triangleright$ Random unit axis of rotation
11:  $\theta_{\max}\sim\text{Beta}(1.0,5.0)\cdot 720^{\circ}\cdot d_{t}$ $\triangleright$ Random angular displacement
12:  Construct cubic spline $s(t)$ on $t_{s},t_{e}]$ with random boundary derivatives
13:  for $f=0$ to F1F-1 do $\triangleright$ Loop over the $F$ frames
14:   $\theta_{f}\leftarrow\text{clamp}(s(f/F),0,1)\,\theta_{\max}$ $\triangleright$ Compute the rotation angle
15:   $\mathbf{R}^{i}_{f}\leftarrow\exp([\theta_{f}\,\mathbf{a}]_{\times})$ $\triangleright$ Compute the rotation matrix
16:  end for
17:end for
18:for $f=0$ to F1F-1 do
19:  $\mathbf{R}_{\text{pano},f}\leftarrow\Pi_{i=1}^{N}\mathbf{R}^{i}_{f}$ $\triangleright$ Combine the NN rotations together
20:end for
21:return $\mathbf{R}_{\text{pano}}$
Prompt AC3D [bahmani2025ac3d] + cam. text. PreciseCam [bernal2025precisecam] + WAN  [wan2025] Ours
A luxurious cliffside resort overlooks a lush landscape, with vibrant greenery and a distant town visible beneath a bright, clear sky, evoking a sense of serene exclusivity. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
A quiet stone church with red pews, arched windows, and historical architecture exudes reverence and solemnity through its timeless design and natural lighting. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Prompt AC3D [bahmani2025ac3d] + cam. text. PreciseCam [bernal2025precisecam] + WAN [wan2025] Ours
A poised wingsuit flyer stands on a grassy mountain peak, overlooking a forested valley and a distant town beneath an overcast sky, evoking tension and grandeur. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
A lively pedestrian street lined with traditional shops and trees, bathed in soft afternoon light, exudes a relaxed, vibrant energy amid a crowd of people. Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative results on the SpatialVID-extreme benchmark. The input absolute camera angle is shown as a dark overlay on the image. We observe that our method aligns well with the overlay, showing the sky when the zenith (“UP”) is in the view and showing the horizon aligned with the equator line (going through “FRONT”), for example. Please refer to the supp. mat. for additional results.
3.3 Null-pitch conditioning
We observed that captioning videos using their actual rotated crops causes the model to overemphasize the textual prompt when determining the camera angle, rather than strictly adhering to the Plücker conditioning. For instance, if a prompt includes ground-level features like grass or asphalt but the camera condition is oriented toward the sky, the model tends to ignore the intended camera control and produce a downward-looking viewpoint. To mitigate this entanglement, we caption each video using an upright, forward-looking image with a 90° field of view, which strengthens the model’s adherence to the camera conditioning (see fig. 5). We refer to this strategy as “null-pitch conditioning”. Specifically, we generate two sets of crops: one using the original randomly sampled camera trajectory, and another with pitch and roll set to 0° and yaw unchanged. As illustrated in fig. 2, a trajectory that consistently points towards the ground would otherwise yield captions describing only ground features. A model trained on such captions would then expect these elements to appear in its outputs, even when the camera conditioning specifies a view of the sky, leading to semantically incorrect generations. Training with null-pitch captions instead provides descriptions of the full scene, making it easier for the model to handle extreme camera angles. During training, the randomly sampled perspective crops are used for the diffusion loss, while the null-pitch crops are used for captioning. We found that applying this scheme to 50% of the samples provides a good balance between camera controllability and prompt adherence, as shown in fig. 5.
4 Experiments
This section presents the training data, provides details on model training and losses, introduces our novel evaluation benchmark, presents baselines, and shows quantitative and qualitative experimental results on our benchmark.
4.1 Training data
We use the WEB360 dataset [wang2024360dvd], which contains geometrically calibrated panoramic videos with leveled horizons. From each 360° frame, we extract six perspective crops and apply ViPE [huang2025vipe] to obtain camera pose annotations, as detailed in Sec. 3.2. The estimated camera poses are then combined with the random crop sampling to obtain the final camera poses, as described in Sec. 3.2.
During training, we sample a camera trajectory using Algorithm 1. We use InternVL-3-2B [zhu2025internvl3] for captioning, providing it with 6 evenly spaced frames across the video timeline. This is also done during training since new videos from different trajectories are sampled each iteration.
Table 1: Quantitative results on our SpatialVID-extreme evaluation dataset. Our method outperforms all baselines on both relative and absolute camera rotation metrics while achieving competitive performance on content alignment (CLIP) and aesthetics (FID, FVD).
Method PitchErr. (abs.) $\downarrow$ GravityErr. (abs.) $\downarrow$ RotErr (rel.) $\downarrow$ TransErr $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FVD $\downarrow$
[…]
4.2 Model and losses
We adopt the same lightweight ControlNet architecture as AC3D [bahmani2025ac3d], using CogVideoX-2B [yang2024cogvideox] as the backbone diffusion model. Following their method, we train the model only on the first 40% of diffusion timesteps and use 8 layers of the original network for the ControlNet training, leaving the remaining layers frozen. We use the same training loss with normalization [karras2024analyzing], and change the relative Plücker-ray representation to our absolute Plücker representation for camera control. We train on four NVIDIA A100 GPUs with 80GB VRAM until convergence ($\sim 8,000$ iterations). See the supplementary material for details and results with our method trained on the WAN [wan2025] backbone.
4.3 Evaluation benchmark
There exists no standard evaluation dataset containing paired text prompts, camera poses, and reference videos that provide extreme camera angles. Most methods [bahmani2025ac3d, he2025cameractrl] use test splits from video datasets like RealEstate10K [RealEstate10K], which have very limited diversity in terms of pitch and roll angles, as shown in fig. 4 and the supplementary material.
We propose a new evaluation benchmark, featuring diverse, high-quality videos, distributed with near-uniform pitch and roll coverage. We start with 371K video clips from the “HQ” subset of the SpatialVID [wang2025spatialvidlargescalevideodataset] dataset, which still exhibit a bias toward forward-looking orientations. We estimate the average pitch from the videos using Perspective Fields [jin2023perspective] and organize them into 10-degree bins spanning -85 to +85 degrees. We manually filter out videos where the pitch estimation clearly failed until we obtain a subset of 140 videos with a uniform pitch distribution. For the roll, we found that even among the 371K videos, nearly all have near-zero roll. We thus artificially add a roll angle by randomly sampling a roll trajectory between -40° and 40° and applying it by warping the video. We name the resulting dataset SpatialVID-extreme. By having a diverse set of pitches and rolls, we ensure a diverse set of text prompts and reference videos for FID and FVD computations. Please refer to the supplementary material for an evaluation of the diversity of our benchmark dataset.
For the camera trajectories fed to the models, we reuse the original camera pose translations, but discard the rotation. We instead generate a random rotation by sampling roll, pitch, and yaw trajectories (see the supp. for details). This strategy allows for a broader range of rotations and also better tests the methods’ adherence to the camera trajectory. Indeed, we receive a normal environment description as input prompt, but for extreme camera angles, such as looking straight up, adhering to the camera conditioning likely requires generating views of the sky.
Table 2: Ablation study on null pitch conditioning. Incorporating null-pitch conditioning improves performance across all metrics and enhances the model’s robustness.
Method PitchErr. (abs.) $\downarrow$ GravityErr. (abs.) $\downarrow$ RotErr (rel.) $\downarrow$ TransErr $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FVD $\downarrow$
w/o null pitch conditioning 28.75 31.70 21.55 0.76 21.29 113.90 983.50
[…]
4.4 Baselines and metrics
PreciseCam + WAN.
A straightforward approach to camera-controlled video generation is to first use a camera-conditioned image generator to synthesize the initial frame, and then apply an image-to-video model to produce the full sequence. This baseline follows that strategy: we use PreciseCam [bernal2025precisecam] to generate the first frame conditioned on the absolute pitch, roll, and field of view, and then use a relative camera control model, WAN-2.1-Fun-I2V Camera Control [wan2025], to extend it into a video.
AC3D + cam. text.
We employ AC3D [bahmani2025ac3d], a text-to-video model with camera control, as baseline. We use their pre-trained checkpoint that uses the same CogVideoX-2B as our model. To provide it with the required initial and final camera poses, we append a text description of the absolute poses, generated from a fixed template, to the original content prompt (refer to the supp. for details).
AC3D + cam. text. + abs. Plücker
This baseline extends the previous one by replacing the relative camera poses with our absolute camera poses. We keep the same model and textual camera descriptions.
Metrics.
To evaluate camera rotation controls, we first measure absolute orientation accuracy using two metrics: angular pitch error, and gravity error, defined as the angle between the ground-truth up vector and the up vector of the generated video, in camera space. For each frame, we extract the pitch and up vector using Perspective Fields [jin2023perspective]. We do not directly evaluate roll error, as estimating it near the poles (looking straight up or down) is unreliable; however, roll deviations are implicitly reflected in the gravity error.
To evaluate the relative rotation error (RotErr), we run VGGT on the generated frames and compute the average relative angular error. We reuse the same definition of rotation error as [he2024cameractrl]. We evaluate content adherence using the CLIP score [hessel2021clipscore] to measure the alignment between the generated video frames and the reference text description. We employ the Fréchet Inception Distance (FID) and Fréchet Video Distance (FVD) to assess the visual quality and realism of the generated frames and sequences, respectively. To evaluate camera motion accuracy, we compute a translation error by comparing the estimated and ground-truth trajectories; both trajectories are normalized prior to computing the error to account for scale differences.
4.5 Results
We present the result of our evaluation in Table 1, which shows that our method clearly outperforms existing methods. Without any change to the architecture, our method nearly halves the pitch and gravity errors and significantly reduces the relative rotation error. We attribute the slight decrease in performance on CLIP, FID and FVD scores to differences in the training data: AC3D is trained on all 65k videos and prompts from RealEstate10K, whereas we only train on 2K 360 videos from WEB360.
We showcase qualitative results in Fig. 6. We observe that while AC3D follows the motion correctly (e.g., the camera progressively looks down throughout the video), it fails to capture the initial target view direction. Feeding it instead the absolute camera poses, which is out of its training distribution, results in negligible improvements in absolute metrics but generates noticeable artefacts (see supp.). PreciseCam [bernal2025precisecam] coupled with WAN-I2V [wan2025] exhibits several failure cases, in particular when used with extreme or unusual camera angles. This is because the input prompt describes the overall environment, which may not be what the camera is pointed at. For example, the prompt in fig. 6 (top left) describes ground elements while the camera aims at the sky. We observe that PreciseCam focuses on text content and often fails to produce the correct viewpoint when it can’t place the content in a semantically meaningful way within the image. In contrast, we use our null-pitch conditioning to ensure that generic descriptions of environments do not prevent the generation of accurate camera viewpoints, as shown in fig. 6 (right).
Ablations.
We ablate our proposed null-pitch conditioning in Tab. 2, which shows that removing it leads to a 20% and 17% drop in absolute camera pitch and gravity accuracy, respectively. This is corroborated qualitatively in fig. 5: in the top and middle rows, where null-pitch conditioning is omitted, the model incorrectly introduces elements such as grass even when the camera is oriented fully upward.
5 Discussion
While our method introduces a novel capability for text-to-video models, it has a few limitations. First, our current on-the-fly camera sampling supports only rotational motion. Extending to translations—potentially by leveraging recent successes in real-time novel view synthesis methods such as Gaussian Splats—remains an interesting direction for future work. Second, the generated videos can exhibit visual artifacts, as visible in the generated crowd in fig. 6. We expect that continued progress in video generation models will mitigate these issues while remaining compatible with our framework.
6 Conclusion
We presented GimbalDiffusion, the first method to enable absolute camera control in text-to-video generation by using gravity as global reference. Our data pipeline combines 360° panoramic videos with camera pose estimation to produce gravity-aligned camera pose annotations. Leveraging panoramic videos allows us to sample the full sphere of possible viewpoints, reducing the domain gap for rare but expressive camera angles. In addition, we proposed null-pitch conditioning, a data generation step that enhances camera conditioning fidelity when the model is given prompts with conflicting cues.
We believe our work lays the foundation for physically grounded camera control in video generation, empowering creators to produce dynamic storyboards and convey their visual ideas with greater precision and speed.
Acknowledgements.
This research was supported by Adobe and a Natural Sciences and Engineering Research Council of Canada (NSERC) scholarship, reference number 600578. Computing resources were provided by Adobe and the Digital Research Alliance of Canada. The authors thank Yohan Poirier-Ginter, Qitao Zhao, Rahul Sajnani, Jack Hilliard and Jonathan Roussel for discussions and proofreading help.

### v2

@@
 Abstract
-Recent progress in text-to-video generation has achieved remarkable realism, yet fine-grained control over camera motion and orientation remains elusive.
-Existing approaches typically encode camera trajectories through relative or ambiguous representations, limiting explicit geometric control.
+Recent progress in text-to-video generation has achieved remarkable realism, yet fine-grained control over camera motion and orientation remains elusive, especially with extreme trajectories (e.g., a 180° turnaround, or looking directly up or down).
+Existing approaches typically encode camera trajectories using relative or ambiguous representations, limiting precise geometric control and offering limited support for large rotations.
 We introduce GimbalDiffusion, a framework that enables camera control grounded in physical-world coordinates, using gravity as a global reference.
-Instead of describing motion relative to previous frames, our method defines camera trajectories in an absolute coordinate system, allowing precise and interpretable control over camera parameters without requiring an initial reference frame.
-We leverage panoramic $360^{\circ}$ videos to construct a wide variety of camera trajectories, well beyond the predominantly straight, forward-facing trajectories seen in conventional video data.
-To further enhance camera guidance, we introduce null-pitch conditioning, an annotation strategy that reduces the model’s reliance on text content when conflicting with camera specifications (e.g., generating grass while the camera points towards the sky).
-Finally, we establish a benchmark for camera-aware video generation by rebalancing SpatialVID-HQ for comprehensive evaluation under wide camera pitch variation.
-Together, these contributions advance the controllability and robustness of text-to-video models, enabling precise, gravity-aligned camera manipulation within generative frameworks.
-“A serene aerial view captures a boat drifting down a dark river, framed by lush greenery and shimmering water under a peaceful sky.”
-Refer to caption Refer to caption Refer to caption
-“A sleek, modern home with high ceilings, natural light, and contemporary furnishings features a man in a blue suit in its open-plan living area.”
-Refer to caption Refer to caption Refer to caption
-Reference Ours AC3D [bahmani2025ac3d]
-Figure 1: Reproducing real scenes using our representation.
-From reference frames from a driving video (left), GimbalDiffusion can be controlled to replicate its camera viewpoint with respect to gravity (center).
-In contrast, existing methods (right) do not accurately match the absolute reference pose.
+Instead of describing motion relative to previous frames, our method defines camera trajectories in an absolute coordinate system, allowing accurate, interpretable control over camera parameters.
+Using panoramic $360^{\circ}$ videos for training, we cover the full sphere of possible viewpoints, including combinations of extreme pitch and roll that are out-of-distribution of conventional video data.
+To improve camera guidance, we introduce null-pitch conditioning, a strategy that prevents the model from overriding camera specifications in the presence of conflicting prompt content (e.g., generating grass while the camera points toward the sky).
+Finally, we propose new benchmarks to evaluate gravity-aware camera-controlled video generation, assessing models’ ability to generate extreme camera angles and quantify their input prompt entanglement.
 1 Introduction
-Text-to-video generation is rapidly advancing, offering astonishing photorealism and new creative workflows.
-Yet, the ability to control generated video content—especially camera movement and trajectory—remains limited.
+Text-to-video generation is rapidly advancing, offering increasingly photorealistic results and new creative workflows.
+Yet, the ability to control generated video content—especially camera trajectory—remains limited, particularly for extreme camera rotations such as top-down shots, Dutch angles, or full 360° pans.
 Early approaches rely solely on textual input which does not provide fine-grained control, such as specifying geometric parameters for camera paths and angles.
-Recent methods [he2024cameractrl, he2025cameractrl, bahmani2025vd3d, bahmani2025ac3d, bai2025recammaster, wan2025] address this limitation by conditioning video models with explicit camera trajectories, frequently encoded as Plücker rays.
+Recent methods [13, 14, 3, 2, 4, 34] address this limitation by conditioning video models with explicit camera trajectories, frequently encoded as Plücker rays.
 However, these representations are inherently ambiguous and lack the precision of real camera systems.
 Typically, camera extrinsics used for Plücker conditioning are defined relative to the first frame, making it impossible to explicitly control camera orientation with an absolute reference within the environment (e.g., with respect to gravity).
+As a result, these methods are limited to modest rotations and fail at extreme camera angles.
 In this work, we introduce GimbalDiffusion, a novel gravity-centric approach to data annotation that overcomes this challenge and enables precise, absolute user control over generated camera trajectories, as illustrated in LABEL:fig:Teaser.
@@
 To resolve rotational ambiguity, we leverage geometrically calibrated 360° panoramic videos.
-By sampling camera crops over the sphere, we can remove the camera bias in human-captured datasets, allowing rare camera shots such as Dutch angles (45° roll) or extreme pitches to be sampled.
-This alignment enables the generation of cinematic shots, such as low- and high-angle perspectives, which are uncommon in regular video datasets.
+By sampling camera crops over the sphere, we can remove camera bias in human-captured datasets, enabling the sampling of rare cinematic camera shots, such as barrel rolls (a “twisting” camera rotation), extreme low- and high-angle perspectives, or looping 360° rotations.
 The effectiveness of this approach is demonstrated in fig.
-1, where our method produces frames with camera views that more closely match the intended trajectories than previous methods.
-To design our dataset from panoramic $360^{\circ}$ videos, original translation paths are preserved while camera rotations are randomly sampled.
-This lets our model learn a wide variety of camera trajectories, far beyond the predominantly straight, forward-facing paths seen in conventional video data.
+1, where our method produces frames whose camera views more closely match the intended trajectories than those of previous methods.
+This enables our model to learn a wide range of camera trajectories, far beyond the predominantly straight, forward-facing paths in conventional videos.
+Forward Up Back Down Forward
+Ours
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+UCPE
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Gen3C
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A spacious living room with a gray sofa facing a wooden coffee table.
+A large window behind sheer curtains lets soft afternoon light enter the room.
+A bookshelf filled with books lines one wall, and a floor lamp stands near the sofa.”
+Figure 1: Extreme rotation handling.
+Here, the camera starts forward and performs a motion akin to a “backflip”: looking straight up, then backwards, down, and back to the starting position.
+Our method (top) precisely follows the desired camera trajectory while accurately reproducing the original content from the first frame.
+Other methods, like UCPE [50] and GEN3C [30] (initialized with PreciseCam [5]), fail at this task.
+Refer to the supp.
+for the full videos.
 We further observe that text and camera angles are inherently entangled, and that care must be taken to ensure that appropriate control is achieved.
-For example, when asked to generate a video of a “green field under a blue sky” with a camera looking down, a model must learn to render grass, even if not specified in the prompt.
-To this end, we propose a null-pitch conditioning approach, where captioning is performed on forward-facing crops rather than the actual camera point of view.
-Finally, we propose a new evaluation benchmark for video generation with extreme camera angles.
-To do so, we use a large-scale dataset of high-quality internet videos, SpatialVID-HQ [wang2025spatialvidlargescalevideodataset], and rebalance the camera pitch to obtain diverse test samples.
-We will release this test dataset upon publication.
+For example, when asked to generate a video of a “green field under a blue sky” with a camera looking down, a model must learn to render only grass, even if other concepts are specified in the prompt.
+To this end, we propose a null-pitch conditioning approach in which captioning is performed on forward-facing crops rather than on the actual camera point of view.
+Finally, we propose two new evaluation benchmarks.
+The first evaluates video generation with extreme camera angles by leveraging the large-scale, high-quality dataset SpatialVID-HQ [35] and rebalancing the camera pitch distribution to obtain diverse test samples.
+The second evaluates entanglement between prompt and camera angle by annotated crops of panoramas from the PolyHaven dataset [12].
+We will release these test datasets upon publication.
 In summary, we make the following contributions:
   • A pipeline for generating diverse camera trajectories from panoramic 360° video, providing exhaustive coverage of possible camera orientations;
-  • A strategy named null-pitch conditioning for disentangling text from absolute camera angle;
-  • A novel benchmark for video generation with extreme camera angles.
-Figure 2: Training data pipeline.
-(a) We extract the camera poses from a dataset of 360° panoramic videos in equirectangular format (see Sec.
-3.2).
-(b) We randomly sample a camera rotation trajectory on the sphere, from which we generate field of view masks in equirectangular format for both the projected image and the null-pitch conditioning (see Sec.
-3.3).
-(c) We integrate the two previous steps by combining the sampled rotation with the estimated camera poses.
-This process creates a dataset of perspective images with their camera poses, along with null-pitch reference images.
-The main advantage of this approach is the ability to sample camera viewing directions across the entire sphere, in contrast to using natural human-captured videos, which are heavily biased towards null roll and pitch.
-Additionally, we generate a text description for both the perspective video and the null-pitch video using a VLM.
+  • A strategy named null-pitch conditioning for disentangling text from absolute camera angle, critical for extreme viewpoints where prompt content and camera orientation diverge most;
+  • Two evaluation benchmarks for: video generation with extreme camera angles, and assessing the entanglement between input prompt and camera orientation.
 2 Related Work
 Video generation.
-Generative video modeling has evolved rapidly, from early GAN-based methods that produced short, low-resolution clips [tulyakov2018mocogan, li2018video, Acharya2018] to large-scale diffusion-based architectures capable of generating temporally coherent, semantically consistent videos [Weng2024].
-Recent open-source models such as CogVideoX [yang2024cogvideox] and WAN [wan2025] push the frontier of video length, resolution, and controllability while proposing both text-to-video and image-to-video variants.
-Editing generated videos remains challenging: precise modification to content, viewpoint, or lighting is difficult once frames are synthesized.
-To address this, several control strategies have emerged, introducing user-guided conditioning based on edge maps, depth, sketches, or explicit 3D geometry for structural control [chen2023controlavideo, Yenphraphai2024ImageSculpting, chang2024magicpose, hu2023animateanyone], and illumination-aware guidance for lighting control [kocsis2024lightit, Magar_lightlab].
-These mechanisms allow precise control over motion, composition, and scene appearance, substantially improving the usability and compositional fidelity of generative video systems.
+Generative video modeling has evolved rapidly, from early GAN-based methods that produced short, low-resolution clips [32, 26, 1] to large-scale diffusion-based architectures capable of generating temporally coherent, semantically consistent videos [40].
+Recent open-source models such as CogVideoX [46] and WAN [34] push the frontier of video length, resolution, and controllability while proposing both text-to-video and image-to-video variants.
+Editing videos remains challenging: precise changes to content, viewpoint, or lighting are difficult once the frames have been synthesized.
+To address this, control strategies have emerged, introducing user-guided conditioning based on edge maps, depth, sketches, or explicit 3D geometry for structural control [7, 47, 6, 16], and illumination-aware guidance for lighting control [23, 28].
+These mechanisms allow precise control over motion, composition, and scene appearance, enhancing the usability and compositional fidelity of generative video systems.
 Camera control methods.
-Beyond content and appearance, viewpoint control has become key in generative video.
-A common representation for such control is Plücker rays, described in Sec.
-3.1.
-This formulation enables differentiable, geometrically consistent camera manipulation.
-In static image generation, PreciseCam [bernal2025precisecam] introduces absolute control over camera pose by conditioning on roll, pitch, field of view, and lens distortion alongside textual prompts.
-For image-to-video generation, methods such as MotionCanvas [xing2025motioncanvas], Diffusion-as-Shader [gu2025das], and TrajectoryCrafter [mark2025trajectorycrafter] employ point cloud conditioning to maintain spatial coherence while navigating a scene from an initial reference image.
-While effective, these approaches depend on the depth and pose estimation quality of the initial reference frame, which can restrict flexibility and generalization.
-In text-to-video synthesis, AC3D [bahmani2025ac3d] allows user-defined camera trajectories but conditions motion relative to the first frame only, leaving global rotational and translational ambiguities unresolved.
+In static image generation, PreciseCam [5] (similarly, [27]) provide text-to-image models with absolute control over camera pose by conditioning on roll, pitch, field of view, and lens distortion alongside textual prompts.
+For image-to-video generation, ViewCrafter [49] and GEN3C [30] generate novel views from a single input image, but cannot handle scene motion.
+Methods such as MotionCanvas [45], Diffusion-as-Shader [11], and TrajectoryCrafter [48] maintain spatial coherence with point cloud conditioning from an initial reference image.
+While effective, these approaches depend on the reference frame’s depth and pose estimation quality, limiting flexibility and generalization.
+In text-to-video synthesis, AC3D [2] allows user-defined camera trajectories but conditions motion relative to the first frame only, leaving global rotational and translational ambiguities unresolved.
 Similarly, Li et al.
-[li2025cameras] manipulate positional encodings to control viewpoint changes, yet still lack a consistent absolute reference such as gravity.
-These works underscore the need for representations that enable precise, absolute camera control and scalable trajectory generation across space and time.
-Our approach addresses this gap by aligning camera orientation with gravity and anchoring motion within a consistent world frame, advancing controllable text-to-video synthesis.
+[25] manipulate positional encodings to control viewpoint changes, yet still lack a consistent absolute reference such as gravity.
+Concurrent with our work, UCPE [50] extends positional encodings to encode a latitude up map.
+However, its data generation is complex and requires matching 360° videos with rotations extracted from a secondary dataset.
+It also cannot handle extreme camera angles (see fig.
+1).
+These works underscore the need for precise, absolute camera control that remains robust under extreme rotations.
+Our approach addresses this gap by aligning camera orientation with gravity and anchoring motion within a consistent world frame, allowing extreme camera rotations in text-to-video synthesis.
+Panorama video generation methods
+[37, 43, 51] adapt video generation backbones for the task of generating 360° equirectangular videos.
+In addition to text, they can be conditioned on an input image [37, 51], optical flow [37] and derotated flow [51].
+While gravity-aligned camera control can be achieved by cropping the desired field of view out of the generated panorama, this comes with several disadvantages: 1) most of the panorama pixels are discarded, resulting in severe resolution degradation; 2) the desired concepts described in the prompt can be cropped out and thus not appear in the video; and 3) current text-conditioned methods [43, 37] do not provide control over camera translation.
+We demonstrate these downsides in the supplementary material.
 Camera pose estimation.
 Estimating accurate camera poses from images is a long-standing problem in 3D vision.
-Classical structure-from-motion pipelines such as VisualSFM [wu2011visualsfm], OpenMVG [moulon2016openmvg], and COLMAP [Schonberger2016COLMAP] reconstruct camera intrinsics and extrinsics through feature matching, triangulation, and global bundle adjustment.
-More recent learning-based approaches jointly infer scene geometry and pose from data: VGGT [wang2025vggt] leverages transformer-based global geometry tokens for pose regression, while DUSt3R [wang2024dust3r] and MASt3R [leroy2024grounding] predict dense correspondence fields to recover metrically scaled 3D structure.
-MegaSAM [Jiang2024MegaSAM] extends these ideas to foundation scale, achieving state-of-the-art alignment accuracy.
+Classical structure-from-motion pipelines such as VisualSFM [42], OpenMVG [29], and COLMAP [31] reconstruct camera intrinsics and extrinsics through feature matching, triangulation, and global bundle adjustment.
+More recent learning-based approaches jointly infer scene geometry and pose from data: VGGT [36] leverages transformer-based global geometry tokens for pose regression, while DUSt3R [39] and MASt3R [24] predict dense correspondence fields to recover metrically scaled 3D structure.
+MegaSAM [18] extends these ideas to foundation scale, achieving state-of-the-art alignment accuracy.
+ViPE [17] extends these ideas to handle in-the-wild videos.
 Despite their accuracy, these systems typically operate in arbitrary coordinate frames and lack physical grounding in gravity or a global world reference, limiting their absolute interpretability.
-To mitigate this, several works estimate the gravity direction from single images [jin2022PerspectiveFields, veicht2024geocalib, xian2019uprightnet].
-However, acquiring datasets with gravity annotations often requires the use of auxiliary sensors, such as accelerometers, or additional semantic information about the scene.
-Examples include EDINA [Do_2022_EgoSceneMSR], Horizon Lines in the Wild [workman2016hlw], and the KITTI Horizon dataset [kluger2020temporally].
-To circumvent such constraints, we leverage geometrically calibrated 360° panoramic videos from WEB360 [huang2025vipe] as backbone for our data generation.
-By combining the up vector derived from the panoramic video with SfM-estimated intrinsics and extrinsics, we obtain a consistent absolute reference for gravity-aware camera control.
-Figure 3: Training data samples from our data augmentation pipeline, capturing a highly diverse set of rotation trajectories from 360° videos.
-Both the trajectory and the prompt are generated on-the-fly in our dataloader.
-Each color in the graph corresponds to a video above.
+To mitigate this, several works estimate the gravity direction from single images [19, 33, 44].
+However, acquiring datasets with gravity annotations often requires auxiliary sensors, such as accelerometers, or additional semantic information about the scene.
+Examples include EDINA [9], Horizon Lines in the Wild [41], and the KITTI Horizon dataset [22].
+ScanNet [8] proposes a dataset of 3D scenes captured using an RGB-D sensor, but exhibits limited diversity, primarily focusing on small indoor environments such as offices, apartments, and bathrooms.
+To circumvent such constraints, we geometrically calibrate 360° videos from the PanoVid dataset [43].
+By combining the up vector derived from the panoramic video with SfM-estimated intrinsics and extrinsics, we obtain a consistent absolute reference for conditioning video models on gravity-aware camera control.
+Figure 2: Training data pipeline.
+On the left, we start by estimating the camera poses and gravity up vectors from a dataset of 360° panoramic videos in equirectangular format (see Sec.
+3.2).
+We then randomly sample a camera rotation trajectory and camera intrinsics on the sphere, which are used for projecting the panoramic frames to perspective images.
+We integrate the two previous steps by combining the sampled rotation with the estimated camera poses (containing also translations).
+This process creates a dataset of perspective images (in green) with their absolute camera poses.
+The main advantage of this approach is the ability to sample camera viewing directions across the entire sphere, in contrast to using natural human-captured videos, which are heavily biased towards null roll and pitch, and minimal rotation amplitude.
+We also project the panorama on the rotation trajectory with pitch and roll removed (in blue), yielding the null-pitch video from which we obtain a textual description (see Sec.
+3.3).
 3 Method
-We aim to train a camera-conditioned model on precise, gravity-referenced camera viewpoints.
+We aim to train a camera-conditioned model that supports extreme camera rotations, including full 360° trajectories, extreme pitch, and large roll, while maintaining precise control.
+To achieve this, we ground camera poses in an absolute, gravity-aligned coordinate system.
 We refer to this global gravity reference as an absolute reference, in contrast to existing relative methods that describe camera motion with respect to a previous frame.
@@
 2.
-In fig.
-2 (a) we extract camera pose annotations from a dataset of 360° panoramic videos, as detailed in Sec.
-3.2.
-Next, in (b), we sample a camera trajectory as described in Sec.
-3.2.
-Finally, in (c), we combine the two previous steps to produce our dataset.
-We repeat the same procedure for an additional version of each trajectory with null pitch and roll, described in Sec.
-3.3.
-Before detailing the full pipeline, the following section explains the standard representation to encode camera poses in neural-network-compatible buffers.
+We first extract camera pose annotations from a dataset of 360° panoramic videos (Sec.
+3.2 and fig.
+2, left).
+Next, we sample a camera trajectory (Sec.
+3.2 and fig.
+2, center-bottom).
+Finally, we combine the two previous steps to produce our dataset.
+We repeat the same procedure for an additional version of each trajectory with null pitch and roll (Sec.
+3.3, fig.
+2, center-top).
+Before detailing the full pipeline, Sec.
+3.1 explains the standard representation to encode camera poses in neural-network-compatible buffers.
 3.1 Representing cameras
-Given a camera extrinsic matrix $\mathbf{E}_{f}\!=\![\mathbf{R}_{f}|\mathbf{t}_{f}]$ and intrinsic matrix $\mathbf{K}_{f}$, a common strategy for conditioning video generation models on viewpoint is to use Plücker rays [kant2024spad, he2024cameractrl, he2025cameractrl, bahmani2025vd3d, bahmani2025ac3d].
+Given a camera extrinsic matrix $\mathbf{E}_{f}\!=\![\mathbf{R}_{f}|\mathbf{t}_{f}]$ and intrinsic matrix $\mathbf{K}_{f}$, a common strategy for conditioning video generation models on viewpoint is to use Plücker rays [20, 13, 14, 3, 2].
 For a conditioning frame $f$, we can define a per-pixel map of Plücker rays as $\mathbf{p}_{f,u,v}\!=\!(\mathbf{t}_{f}\times\mathbf{d}_{f,u,v}^{\prime},\mathbf{d}_{f,u,v}^{\prime})$, where $\mathbf{t}_{f}$ denotes the camera origin in world coordinates, and $\mathbf{d}_{f,u,v}$ is the direction vector from the camera center toward the center of pixel $(u,v)$.
@@
 3.2 Creating varied camera paths from 360° videos
-Most video datasets exhibit a significant bias towards forward-looking views, where the horizon remains level or close to the center of the image, as shown in fig.
-4.
-However, such “normal” viewpoints restrict the expressive range that a filmmaker might wish to achieve in a storyboard.
-For instance, these datasets often lack cinematic shots, such as top-down, low-angle, or Dutch-angle shots.
-As a result, generative video models trained on such data struggle to produce these types of cinematic shots, since they fall outside the training distribution.
-To address this limitation, we propose generating synthetic camera trajectories by sampling from 360° videos and rectifying them so that the horizon position is known throughout each sequence.
+Most video datasets exhibit a significant bias toward forward-looking views with horizon level or near the center of the image, as shown in the supp.
+material.
+However, such “normal” viewpoints limit storytelling expressivity by lacking cinematic shots, such as top-down, low-angle, or barrel-roll shots.
+Consequently, generative video models trained on such data struggle to produce these types of cinematic shots, as they lie outside the training distribution (see fig.
+1).
+To address this limitation, we propose generating synthetic camera trajectories by sampling and rectifying 360° videos.
 Camera poses from SfM.
-From a 360° video, we compute the relative motion between each frame by extracting six perspective crops (front, back, left, right, top, bottom), and processing the resulting video with structure-from-motion (SfM).
-This results in a set of camera poses $\mathbf{E}_{\text{SfM},f}$ at each frame $f$.
+From a 360° video, we compute the relative motion between frames by extracting six perspective crops (front, back, left, right, top, bottom), and processing the resulting video with a pose estimation method [17].
+This yields a set of camera poses $\mathbf{E}_{\text{SfM},f}$ for each frame $f$.
 Taking the first frame of the sequence $f=0$ as reference, we can then compute the relative motion
 $\mathbf{E}_{\text{rel},f}$ $=\mathbf{E}_{\text{SfM},0}^{-1}\mathbf{E}_{\text{SfM},f}\;.$ (1)
+Gravity annotation.
+Pose estimation methods such as [17] provide relative estimates of camera rotations.
+To anchor the videos within a shared global reference frame, we align the relative camera rotations with gravity.
+Specifically, we use [19] to estimate the camera roll and pitch angles at frame 0.
+Because it operates on perspective images, we first project the panoramic frame into eight perspective views evenly sampled about the yaw axis at fixed 90° field of view.
+We then estimate the camera up vectors from these views and rotate them back to the null yaw reference frame.
+We then average the up vectors across these views, yielding $\mathbf{\tilde{u}}_{0}$.
+From this, we derive the absolute, gravity-aligned camera pose for each frame as
+$\mathbf{t}_{\text{abs},f}=\psi(\mathbf{\tilde{u}}_{0})\mathbf{t}_{\text{rel},f}$
+$\mathbf{R}_{\text{abs},f}=\mathbf{R}_{\text{pano},f}\psi(\mathbf{\tilde{u}}_{0})\mathbf{R}_{\text{rel},f}$ (2)
+$\mathbf{E}_{\text{abs},f}=\varphi(\mathbf{R}_{\text{pano},0})[\mathbf{R}_{\text{abs},f}|\mathbf{t}_{\text{abs},f}]\;,$
+where $\varphi(\cdot)$ is a function that preserves only the roll and pitch angles (removes the yaw) of the first sampled camera pose, and $\psi(\cdot)$ aligns camera poses to the gravity up vector.
+After this adjustment, the transformed pose $\mathbf{E}_{\text{abs},f}$ has no rotation about the global up-vector axis.
+Please refer to the supplementary material for the full definition of $\varphi(\cdot)$ and $\psi(\cdot)$.
+This process establishes a gravity-aligned coordinate system consistent across all videos, enabling rotations to be defined within a common global frame.
+Consequently, each sequence is initialized with null yaw, reducing ambiguity and facilitating training with this representation.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“[…] an aerial view of a large resort complex featuring multiple buildings with red roofs […]”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“[…] large outdoor gathering of people and cars. […] The sky is overcast […]”
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A young man in a black hoodie and pants walks along a […] street, engrossed in his smartphone. […]”
+Videos and captions Camera poses
+Figure 3: Training data samples from our data augmentation pipeline, capturing a highly diverse set of rotation trajectories from 360° videos.
+Both the trajectory and the prompt are generated on-the-fly in our dataloader.
+See the supp.
+for full videos.
 Sampling random camera paths.
@@
 We first sample random pitch, roll, and yaw angles uniformly to obtain an initial viewpoint.
-Next, 1 to 4 additional random rotations are sampled and distributed across the sequence length.
+Next, 1 to 3 additional random rotations are sampled and distributed across the sequence length.
 Those rotations are applied sequentially to produce smooth and varied motion paths.
@@
 The FoV values are then interpolated using cubic splines with randomized derivative boundary conditions.
+This approach provides much broader coverage of possible viewpoints than existing datasets (see supplementary).
+Notably, this sampling slightly biases viewpoints towards the poles, which we found to better support model stability and representation of more extreme viewpoints.
 Examples from our sampling are shown in fig.
 3.
-This approach provides broad coverage of the possible viewpoints, as shown in fig.
-4.
-Notably, this sampling slightly biases viewpoints towards the poles, which we found to better support model stability and representation of more extreme viewpoints.
-Gravity annotation.
-To anchor the videos within a shared global reference frame, we align their rotation with gravity. We are able to compute the absolute, gravity-aligned camera pose for each frame as
-$\mathbf{E}_{\text{abs},f}=\varphi(\mathbf{R}_{\text{pano},0})\mathbf{R}_{\text{pano},f}\mathbf{E}_{\text{rel},f}\;,$ (2)
-where $\varphi(\cdot)$ is a function that preserves only the roll and pitch angles (removes the yaw) of the first sampled camera pose.
-After this adjustment, the transformed pose $\mathbf{E}_{\text{abs},f}$ has no rotation about the global up-vector axis.
-This function $\varphi(\cdot)$ is obtained by computing the gravity direction in camera coordinates and inverting it to recover the corresponding up-vector, see the supplementary for the complete definition.
-This process establishes a gravity-aligned coordinate system that is consistent across all videos, allowing rotations to be defined within a common global frame.
-Consequently, each sequence is initialized at null yaw, reducing ambiguity and easing training on this representation.
-Figure 4: Comparison of parameter distribution between our sampling and a typical video dataset for camera control (RealEstate10K [RealEstate10K]).
-Our sampling produces videos with more rotational motion, as measured by total angular distance (the sum of absolute angular differences between consecutive camera poses), and greater diversity in Euler angles (pitch, roll, yaw).
-All values on the $x$-axis are expressed in degrees.
-Up (+90°) Forward (0°) Down (-90°)
-PreciseCam
-Refer to caption Refer to caption Refer to caption
-w/o null pitch
-Refer to caption Refer to caption Refer to caption
-Ours
-Refer to caption Refer to caption Refer to caption
-Prompt: A prairie with grass and flowers, and a beautiful blue sky with few clouds.
-Figure 5: Entanglement between prompt and camera pitch.
-Without a careful captioning strategy, diffusion models may ignore the camera conditioning when the prompt semantics conflict with it.
-Our null-pitch conditioning mitigates this issue and preserves accurate camera control.
-Algorithm 1 Random camera path sampling
-[…]
-[…]
-[…]
-[…]
-[…]
-6:for $i=1$ to NN do
-7:  $d_{t}\sim\text{Uniform}(0,1)$ $\triangleright$ Random rotation duration
-8:  $t_{s}\sim\text{Uniform}(0,1-d_{t})$ $\triangleright$ Random start time
-9:  $t_{e}\leftarrow t_{s}+d_{t}$ $\triangleright$ End time
-10:  $\mathbf{a}\sim\mathbb{S}^{2}$ $\triangleright$ Random unit axis of rotation
-11:  $\theta_{\max}\sim\text{Beta}(1.0,5.0)\cdot 720^{\circ}\cdot d_{t}$ $\triangleright$ Random angular displacement
-12:  Construct cubic spline $s(t)$ on $t_{s},t_{e}]$ with random boundary derivatives
-13:  for $f=0$ to F1F-1 do $\triangleright$ Loop over the $F$ frames
-14:   $\theta_{f}\leftarrow\text{clamp}(s(f/F),0,1)\,\theta_{\max}$ $\triangleright$ Compute the rotation angle
-15:   $\mathbf{R}^{i}_{f}\leftarrow\exp([\theta_{f}\,\mathbf{a}]_{\times})$ $\triangleright$ Compute the rotation matrix
-16:  end for
-17:end for
-18:for $f=0$ to F1F-1 do
-19:  $\mathbf{R}_{\text{pano},f}\leftarrow\Pi_{i=1}^{N}\mathbf{R}^{i}_{f}$ $\triangleright$ Combine the NN rotations together
-20:end for
-21:return $\mathbf{R}_{\text{pano}}$
-Prompt AC3D [bahmani2025ac3d] + cam. text. PreciseCam [bernal2025precisecam] + WAN  [wan2025] Ours
-A luxurious cliffside resort overlooks a lush landscape, with vibrant greenery and a distant town visible beneath a bright, clear sky, evoking a sense of serene exclusivity.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-A quiet stone church with red pews, arched windows, and historical architecture exudes reverence and solemnity through its timeless design and natural lighting.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Prompt AC3D [bahmani2025ac3d] + cam. text. PreciseCam [bernal2025precisecam] + WAN [wan2025] Ours
-A poised wingsuit flyer stands on a grassy mountain peak, overlooking a forested valley and a distant town beneath an overcast sky, evoking tension and grandeur.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-A lively pedestrian street lined with traditional shops and trees, bathed in soft afternoon light, exudes a relaxed, vibrant energy amid a crowd of people.
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
+Algorithm 1 Random camera rotation path sampling
+$\text{pitch}_{0}\sim\text{Uniform}(-90\degree{},90\degree{})$ ; $\text{roll}_{0}\sim\text{Uniform}(-90\degree{},90\degree{})$ ; $\text{yaw}_{0}\sim\text{Uniform}(0\degree{},360\degree{})$
+$\mathbf{R}_{\text{pano},f}\leftarrow\left(\mathcal{R}_{\text{YXZ}}(\text{yaw}_{0},\text{pitch}_{0},\text{roll}_{0})\right)_{f=0}^{F-1}$ $\triangleright$ Initialize the camera orientation
+$N\sim\{0,1,2,3\}$ $\triangleright$ Random number of rotation axis
+for $i=1$ to NN do
+  $d_{t}\sim\text{Uniform}(0,1)$ $\triangleright$ Random rotation duration
+  $t_{s}\sim\text{Uniform}(0,1-d_{t})$ ; $t_{e}\leftarrow t_{s}+d_{t}$ $\triangleright$ Random start and end time
+  $\mathbf{a}\sim\mathbb{S}^{2}$ $\triangleright$ Random unit axis of rotation
+  $\theta_{\max}\sim\text{Beta}(1.0,5.0)\cdot 720^{\circ}\cdot d_{t}$ $\triangleright$ Random angular displacement
+  Construct cubic spline $s(t)$ on $t_{s},t_{e}]$ with random boundary derivatives
+  for $f=0$ to F1F-1 do $\triangleright$ Loop over the $F$ frames
+   $\theta_{f}\leftarrow\text{clamp}(s(f/F),0,1)\,\theta_{\max}$ $\triangleright$ Compute the rotation angle
+   $\mathbf{R}_{\text{pano},f}\leftarrow\exp([\theta_{f}\,\mathbf{a}]_{\times})\times\mathbf{R}_{\text{pano},f}$ $\triangleright$ Update the camera orientation
+  end for
+end for
+return $\mathbf{R}_{\text{pano}}$
+UCPE PreciseCam+GEN3C AC3D+cam. text. Ours UCPE PreciseCam+GEN3C AC3D+cam. text. Ours
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+A sleek, modern interior features a dark, hexagonally lit ceiling that radiates brightness and sophistication, creating a visually striking and inviting atmosphere.
+A modern skyscraper rises against a clear blue sky, its grid-like facade reflecting sunlight in a bustling urban setting.
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+A serene forest is bisected by a powerful waterfall, its vivid greens, yellows, and reds evoking tranquility and natural majesty.
+A sleek, modern home with high ceilings, natural light, and contemporary furnishings features a man in a blue suit in its open-plan living area.
 Figure 4: Qualitative results on the SpatialVID-extreme benchmark.
@@
 Please refer to the supp.
-mat.
 for additional results.
 3.3 Null-pitch conditioning
-We observed that captioning videos using their actual rotated crops causes the model to overemphasize the textual prompt when determining the camera angle, rather than strictly adhering to the Plücker conditioning.
-For instance, if a prompt includes ground-level features like grass or asphalt but the camera condition is oriented toward the sky, the model tends to ignore the intended camera control and produce a downward-looking viewpoint.
-To mitigate this entanglement, we caption each video using an upright, forward-looking image with a 90° field of view, which strengthens the model’s adherence to the camera conditioning (see fig.
-5).
+We observed that captioning videos using their actual rotated crops causes the model to overemphasize the textual prompt when inferring the camera angles, rather than strictly adhering to the Plücker conditioning.
+This entanglement hinders the achievement of extreme camera angles, where visual content diverges most from the prompt.
+For instance, if a prompt describes ground features such as grass but the camera should face the sky, the model often still outputs a downward-looking view.
+To mitigate this entanglement, we caption each video using an upright, forward-looking image with a 90° field of view, thereby strengthening the model’s adherence to the camera conditioning.
 We refer to this strategy as “null-pitch conditioning”.
-Specifically, we generate two sets of crops: one using the original randomly sampled camera trajectory, and another with pitch and roll set to 0° and yaw unchanged.
+Specifically, we generate two sets of crops (see fig.
+2): one using the original randomly sampled camera trajectory, and another with pitch and roll set to 0° (with yaw unchanged).
 As illustrated in fig.
@@
 A model trained on such captions would then expect these elements to appear in its outputs, even when the camera conditioning specifies a view of the sky, leading to semantically incorrect generations.
-Training with null-pitch captions instead provides descriptions of the full scene, making it easier for the model to handle extreme camera angles.
+Training with null-pitch captions, which provide descriptions of the full scene, makes it easier for the model to handle extreme camera angles.
 During training, the randomly sampled perspective crops are used for the diffusion loss, while the null-pitch crops are used for captioning.
-We found that applying this scheme to 50% of the samples provides a good balance between camera controllability and prompt adherence, as shown in fig.
-5.
+While this strategy helps the video model in handling extreme angles, we found it still struggled with undesirable elements present directly below the camera in real 360° training videos, such as tripods or the hand of a person holding a selfie stick.
+Indeed, since these elements are never present in the training caption, the model learns to generate these objects as if they were normal scene content.
+To prevent this, we generate a third set of videos (not shown in fig.
+2), this time looking straight down (-90° pitch), and caption them with the VLM.
+We append the resulting “look-down” caption to the null-pitch caption 50% of the time.
+At inference time, this “look-down” prompt is used in the classifier-free guidance as a negative prompt to ensure that undesirable content is absent from the generated videos.
+Please see the supplementary for additional details.
 4 Experiments
-This section presents the training data, provides details on model training and losses, introduces our novel evaluation benchmark, presents baselines, and shows quantitative and qualitative experimental results on our benchmark.
+This section presents the training data, details model training and losses, introduces our two novel evaluation benchmarks, presents baselines, and shows quantitative and qualitative experimental results on our benchmarks.
 4.1 Training data
-We use the WEB360 dataset [wang2024360dvd], which contains geometrically calibrated panoramic videos with leveled horizons.
-From each 360° frame, we extract six perspective crops and apply ViPE [huang2025vipe] to obtain camera pose annotations, as detailed in Sec.
+We use the “YouTube videos” subset of the PanoVid dataset [43], which contains a diverse set of high-quality 360° videos, and filter out videos where the shortest side is less than 900 pixels.
+We obtain relative camera poses and per-frame up-vectors using the procedure from Sec.
 3.2.
-The estimated camera poses are then combined with the random crop sampling to obtain the final camera poses, as described in Sec.
-3.2.
-During training, we sample a camera trajectory using Algorithm 1.
-We use InternVL-3-2B [zhu2025internvl3] for captioning, providing it with 6 evenly spaced frames across the video timeline.
+We discard videos where pose estimation either failed or resulted in extreme mean acceleration values, typically indicating unreliable camera poses.
+Our filtering strategy reduces the number of training videos from 7797 to 4798 with high-quality camera poses (38% discarded).
+During training, we sample a random camera trajectory using Algorithm 1 at a resolution of $720\times 480$, effectively acting as an infinite data augmentation strategy.
+We use InternVL-3-2B [52] for captioning, providing it with 6 evenly spaced frames across the video timeline.
 This is also done during training since new videos from different trajectories are sampled each iteration.
-Table 1: Quantitative results on our SpatialVID-extreme evaluation dataset.
-Our method outperforms all baselines on both relative and absolute camera rotation metrics while achieving competitive performance on content alignment (CLIP) and aesthetics (FID, FVD).
-Method PitchErr. (abs.) $\downarrow$ GravityErr. (abs.) $\downarrow$ RotErr (rel.) $\downarrow$ TransErr $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FVD $\downarrow$
-[…]
-[…]
-[…]
-[…]
+Prompt lengths are further sampled between short, medium and long.
 4.2 Model and losses
-We adopt the same lightweight ControlNet architecture as AC3D [bahmani2025ac3d], using CogVideoX-2B [yang2024cogvideox] as the backbone diffusion model.
-Following their method, we train the model only on the first 40% of diffusion timesteps and use 8 layers of the original network for the ControlNet training, leaving the remaining layers frozen.
-We use the same training loss with normalization [karras2024analyzing], and change the relative Plücker-ray representation to our absolute Plücker representation for camera control.
-We train on four NVIDIA A100 GPUs with 80GB VRAM until convergence ($\sim 8,000$ iterations).
-See the supplementary material for details and results with our method trained on the WAN [wan2025] backbone.
-4.3 Evaluation benchmark
+We start from the WAN-2.2 5B text-to-video model [34], and add a lightweight camera encoder consisting of two convolution layers and a group normalization layer (see supp.
+for details).
+The last convolution is zero-initialized to avoid instabilities at the beginning of training.
+Its output is then added after initial patchify layer of the DiT backbone.
+During training, both the camera encoder and the backbone are finetuned for 30k iterations (1 day on 8 NVIDIA H200 GPUs).
+We use the AdamW optimizer with a learning rate of $10^{-5}$ and a weight decay of 0.01.
+4.3 Video camera control benchmark
 There exists no standard evaluation dataset containing paired text prompts, camera poses, and reference videos that provide extreme camera angles.
-Most methods [bahmani2025ac3d, he2025cameractrl] use test splits from video datasets like RealEstate10K [RealEstate10K], which have very limited diversity in terms of pitch and roll angles, as shown in fig.
-4 and the supplementary material.
+Most methods [2, 14] use test splits from video datasets like RealEstate10K [10], which have very limited diversity in terms of pitch and roll angles, as shown in the supplementary material.
 We propose a new evaluation benchmark, featuring diverse, high-quality videos, distributed with near-uniform pitch and roll coverage.
-We start with 371K video clips from the “HQ” subset of the SpatialVID [wang2025spatialvidlargescalevideodataset] dataset, which still exhibit a bias toward forward-looking orientations.
-We estimate the average pitch from the videos using Perspective Fields [jin2023perspective] and organize them into 10-degree bins spanning -85 to +85 degrees.
+We start with 371K video clips from the “HQ” subset of the SpatialVID [35] dataset, which still exhibit a bias toward forward-looking orientations.
+We estimate the average pitch from the videos using Perspective Fields [19] and organize them into 10-degree bins spanning -85° to +85°.
 We manually filter out videos where the pitch estimation clearly failed until we obtain a subset of 140 videos with a uniform pitch distribution.
@@
 Please refer to the supplementary material for an evaluation of the diversity of our benchmark dataset.
-For the camera trajectories fed to the models, we reuse the original camera pose translations, but discard the rotation.
+For the camera trajectories fed to the models, we annotate translations in metric scale using ViPE [17], but discard the rotation.
 We instead generate a random rotation by sampling roll, pitch, and yaw trajectories (see the supp.
@@
 Indeed, we receive a normal environment description as input prompt, but for extreme camera angles, such as looking straight up, adhering to the camera conditioning likely requires generating views of the sky.
-Table 2: Ablation study on null pitch conditioning. Incorporating null-pitch conditioning improves performance across all metrics and enhances the model’s robustness.
-Method PitchErr. (abs.) $\downarrow$ GravityErr. (abs.) $\downarrow$ RotErr (rel.) $\downarrow$ TransErr $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ FVD $\downarrow$
-w/o null pitch conditioning 28.75 31.70 21.55 0.76 21.29 113.90 983.50
-[…]
 4.4 Baselines and metrics
+UCPE
+[50] is a concurrent method that also generates videos in a gravity-aligned reference frame, using WAN-2.1 as the backbone.
+Unlike ours, UCPE requires a limited auxiliary dataset to simulate rotation diversity within panoramas, resulting in a less diverse dataset than our proposed approach.
+AC3D+cam. text.
+We employ AC3D [2], a text-to-video model with camera control, as baseline.
+We use their pre-trained checkpoint using CogVideoX-2B as the base video model.
+To provide it with the required initial and final camera poses, we append a text description of the absolute poses, generated from a fixed template, to the original content prompt (refer to the supp.
+for details).
+AC3D+cam. text.+abs. Plücker.
+This baseline extends the previous one by replacing the relative camera poses with our absolute camera poses. We keep the same model and textual camera descriptions.
 PreciseCam+WAN.
 A straightforward approach to camera-controlled video generation is to first use a camera-conditioned image generator to synthesize the initial frame, and then apply an image-to-video model to produce the full sequence.
-This baseline follows that strategy: we use PreciseCam [bernal2025precisecam] to generate the first frame conditioned on the absolute pitch, roll, and field of view, and then use a relative camera control model, WAN-2.1-Fun-I2V Camera Control [wan2025], to extend it into a video.
-AC3D + cam. text.
-We employ AC3D [bahmani2025ac3d], a text-to-video model with camera control, as baseline.
-We use their pre-trained checkpoint that uses the same CogVideoX-2B as our model.
-To provide it with the required initial and final camera poses, we append a text description of the absolute poses, generated from a fixed template, to the original content prompt (refer to the supp.
-for details).
-AC3D + cam. text. + abs. Plücker
-This baseline extends the previous one by replacing the relative camera poses with our absolute camera poses. We keep the same model and textual camera descriptions.
+This baseline follows that strategy: PreciseCam [5] generates the first frame conditioned on the absolute pitch, roll, and field of view, and then WAN-2.1-Fun-I2V Camera Control [34] prolongs that first frame to a video.
+PreciseCam+GEN3C.
+GEN3C [30] takes as input a 3D point cloud from a single image or an input video which can then be rendered from novel camera views.
+We reuse the original implementation which predicts the depth using MoGE [38] and feed it the first frame from PreciseCam.
+We then project the point cloud along the relative camera poses and feed them to GEN3C.
+Note however that GEN3C can only generate novel views of static scenes, and therefore cannot produce dynamic video content (e.g., walking people).
 Metrics.
-To evaluate camera rotation controls, we first measure absolute orientation accuracy using two metrics: angular pitch error, and gravity error, defined as the angle between the ground-truth up vector and the up vector of the generated video, in camera space.
-For each frame, we extract the pitch and up vector using Perspective Fields [jin2023perspective].
-We do not directly evaluate roll error, as estimating it near the poles (looking straight up or down) is unreliable; however, roll deviations are implicitly reflected in the gravity error.
-To evaluate the relative rotation error (RotErr), we run VGGT on the generated frames and compute the average relative angular error.
-We reuse the same definition of rotation error as [he2024cameractrl].
-We evaluate content adherence using the CLIP score [hessel2021clipscore] to measure the alignment between the generated video frames and the reference text description.
-We employ the Fréchet Inception Distance (FID) and Fréchet Video Distance (FVD) to assess the visual quality and realism of the generated frames and sequences, respectively.
-To evaluate camera motion accuracy, we compute a translation error by comparing the estimated and ground-truth trajectories; both trajectories are normalized prior to computing the error to account for scale differences.
+To evaluate camera rotation controls, we first measure absolute orientation accuracy using two metrics: angular pitch error (“PitchErr”), and gravity error (“GravityErr”), defined as the angle between the ground-truth up vector and that of the generated video, in camera space.
+The pitch angle and up vector are obtained from each frame using Perspective Fields [19].
+Roll error is not directly computed, as estimating it near the poles (looking straight up or down) is unreliable; however, roll deviations are implicitly reflected in the gravity error.
+To evaluate the relative rotation error (“RotErr”), we run VGGT [36] on the generated frames and compute the average relative angular error.
+We reuse the same definition of rotation error as [13].
+To evaluate camera motion accuracy, we compute a translation error (“TransErr”) by comparing the estimated and ground-truth trajectories; both trajectories are normalized prior to computing the error to account for scale differences.
+The CLIP score [15] is computed to measure the alignment between the generated video frames and the reference text description.
+Finally, we employ the Fréchet Inception Distance (FID) and Fréchet Video Distance (FVD) to assess the visual quality and realism of the generated frames and sequences, respectively.
+Table 1: Quantitative results on our SpatialVID-extreme evaluation dataset.
+Our method outperforms all baselines on absolute camera orientation metrics (PitchErr and GravityErr), and all dynamic video generation methods on relative rotation error (RotErr), while achieving competitive performance on translation error (TransErr), and aesthetics (FID, FVD).
+Furthermore, removing the null-pitch conditioning drastically degrades absolute orientation metrics.
+Method PitchErr GravityErr RotErr TransErr CLIP FID FVD
+Novel view synthesis methods
+PreciseCam+GEN3C 26.82 31.09 10.99 0.40 20.17 101.41 830.7
+Dynamic video generation methods
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 4.5 Results
-We present the result of our evaluation in Table 1, which shows that our method clearly outperforms existing methods.
-Without any change to the architecture, our method nearly halves the pitch and gravity errors and significantly reduces the relative rotation error.
-We attribute the slight decrease in performance on CLIP, FID and FVD scores to differences in the training data: AC3D is trained on all 65k videos and prompts from RealEstate10K, whereas we only train on 2K 360 videos from WEB360.
-We showcase qualitative results in Fig.
-6.
-We observe that while AC3D follows the motion correctly (e.g., the camera progressively looks down throughout the video), it fails to capture the initial target view direction.
-Feeding it instead the absolute camera poses, which is out of its training distribution, results in negligible improvements in absolute metrics but generates noticeable artefacts (see supp.).
-PreciseCam [bernal2025precisecam] coupled with WAN-I2V [wan2025] exhibits several failure cases, in particular when used with extreme or unusual camera angles.
-This is because the input prompt describes the overall environment, which may not be what the camera is pointed at.
-For example, the prompt in fig.
-6 (top left) describes ground elements while the camera aims at the sky.
-We observe that PreciseCam focuses on text content and often fails to produce the correct viewpoint when it can’t place the content in a semantically meaningful way within the image.
-In contrast, we use our null-pitch conditioning to ensure that generic descriptions of environments do not prevent the generation of accurate camera viewpoints, as shown in fig.
-6 (right).
+We present the quantitative results of our evaluation in Sec.
+4.4.
+Overall, our method outperforms all dynamic video generation baselines on all camera orientation metrics (“PitchErr”, “GravityErr”, and “RotErr”).
+Since GEN3C can only synthesize novel viewpoints without scene motion, we include its results here for reference only.
+Our method achieves FID scores similar to those of other baselines and FVD scores on par with UCPE and AC3D.
+Baselines using PreciseCam as a starting point improve the FVD score, at the cost of much worse camera orientation metrics.
+We observe that our method results in a slight drop in CLIP score, see the discussion on prompt-camera entanglement in Sec.
+4.6 below.
+Representative qualitative results are illustrated in fig.
+4 (see the supplementary for more).
+We observe that all baselines struggle with extreme angles (e.g., the “up” starting point in the bottom-left example).
+Here, they attempt to generate visual content in accordance with the prompt, but this contradicts the desired camera orientation.
+Another example is shown in the top-right example, where the “skyscraper rising against a clear blue sky” eventually becomes visible in “Ours” (as it should when the camera starts to look up), whereas both UCPE and “AC3D+cam.
+text” are incorrectly rendering it right from the first frame.
+In this example, PreciseCam fails to properly generate the first frame.
+Overall, visual examples confirm that our approach provides the best camera control, even in the case of extreme viewpoints.
 Ablations.
-We ablate our proposed null-pitch conditioning in Tab.
-2, which shows that removing it leads to a 20% and 17% drop in absolute camera pitch and gravity accuracy, respectively.
-This is corroborated qualitatively in fig.
-5: in the top and middle rows, where null-pitch conditioning is omitted, the model incorrectly introduces elements such as grass even when the camera is oriented fully upward.
+We ablate our proposed null-pitch conditioning in Sec.
+4.4, which shows that its absence leads to a 67% and 52% increase in camera pitch and gravity error, respectively.
+Next, we investigate the effect of null-pitch conditioning.
+4.6 Prompt-camera entanglement evaluation
+To investigate the effect of null-pitch conditioning, we develop a second benchmark dataset to evaluate whether methods can disambiguate between two possibly conflicting inputs: the prompt and the camera pitch angle.
+Indeed, pitch is highly correlated with semantic concepts, e.g., a +90° pitch (looking up) is strongly associated with images of the sky/ceiling, whereas a -90° pitch is strongly associated with ground features.
+Here, roll and yaw angles are ignored, since both are uncorrelated to object semantics in the field of view.
+We start by randomly selecting 20 panoramas from the PolyHaven dataset [12].
+For each, we take a forward-looking crop with a 90° FoV and caption it using InternVL-3-8B [52].
+We then give this prompt to each evaluated method along with a static camera orientation.
+The methods are evaluated on pitch angles going from -90° to +90°, in 10° increments.
+In total, this benchmark consists of 380 pairs of input prompts and pitch angles.
+The results obtained with baselines providing gravity-aligned camera control (UCPE [50] and PreciseCam [5]) are shown in fig.
+5.
+Qualitatively, we observe in fig.
+5-(a) that our proposed null-pitch conditioning enforces strict adherence to the camera orientation when the camera is pointing straight up or straight down.
+Quantitatively, we observe in fig.
+5-(b) (top) a substantially lower pitch-angle error across all 20 scenes, particularly at extreme absolute pitch values.
+One simple way to assess the visual alignment between the text prompt and the output video is to compute the CLIP score directly between them.
+However, this naive approach has a key limitation: the prompt features may rightly be absent from the video if they violate the user-specified camera constraint (e.g., textual descriptions of ground content should not be visible if the camera is pointing up)—but not doing so would negatively affect the CLIP score.
+To address this contradiction, we compute the CLIP score between the generated video frames and a caption of the ground-truth frame (obtained with InternVL-3-8B) for each pitch angle separately, and report results in fig.
+5-(b) (bottom).
+We observe that null-pitch conditioning improves CLIP similarity at extreme pitch values, aligning results closer to the ground-truth curve.
+See the supplementary material for a detailed analysis.
+PreciseCam UCPE w/o null pitch Ours Ground truth
+Up (+90°)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Forward (0°)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Down (-90°)
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+“A serene landscape unfolds with dry grass swaying gently in the foreground.
+The sun shines brightly in the clear blue sky, casting a warm glow over the scene.
+Sparse green bushes dot the landscape, with distant hills visible on the horizon.
+The scene remains tranquil, with no visible movement or change in the environment.”
+(a) Qualitative example from our entanglement benchmark
+Refer to captionRefer to caption
+(b) Pitch error and CLIP variation over our entanglement benchmark.
+Figure 5: Evaluating the entanglement between prompt and camera pitch.
+(a) Without a careful captioning strategy, generative models may ignore the camera conditioning when it conflicts with the prompt semantics, which is especially problematic at extreme pitch values (e.g., -90° and +90°).
+(b) Quantitative evaluation of this effect: our null-pitch conditioning strategy reduces the aggregated pitch error at extreme angles (top) and maximizes the content similarity with the ground truth, as reflected in the CLIP scores (bottom).
+Here, the CLIP scores are shifted to each method’s average to highlight their variations with respect to the input pitch.
+Refer to the supp.
+for additional results.
 5 Discussion
-While our method introduces a novel capability for text-to-video models, it has a few limitations.
-First, our current on-the-fly camera sampling supports only rotational motion.
-Extending to translations—potentially by leveraging recent successes in real-time novel view synthesis methods such as Gaussian Splats—remains an interesting direction for future work.
-Second, the generated videos can exhibit visual artifacts, as visible in the generated crowd in fig.
-6.
+We present GimbalDiffusion, a method that enables extreme and arbitrary camera control in text-to-video generation by grounding camera poses in a gravity-aligned absolute coordinate system.
+Our pipeline combines 360° panoramic videos with camera pose estimation to produce training data covering the full sphere of possible viewpoints, including extreme pitch, large roll, and full 360° rotation trajectories that were previously beyond the reach of generative video models.
+We also proposed null-pitch conditioning, a data generation step that enhances camera conditioning fidelity when the text prompts contain conflicting cues.
+While it introduces novel capabilities for text-to-video models, our method still has a few limitations.
+First, our current on-the-fly camera sampling is restricted to rotations and must use the translations from the original source videos.
+A promising area for future research is to extend this data augmentation to include translations, potentially by leveraging recent successes in real-time novel view synthesis methods such as 3DGS [21].
+Second, the generated videos can exhibit visual artifacts (for example, the faceless man in fig.
+4).
 We expect that continued progress in video generation models will mitigate these issues while remaining compatible with our framework.
-6 Conclusion
-We presented GimbalDiffusion, the first method to enable absolute camera control in text-to-video generation by using gravity as global reference.
-Our data pipeline combines 360° panoramic videos with camera pose estimation to produce gravity-aligned camera pose annotations.
-Leveraging panoramic videos allows us to sample the full sphere of possible viewpoints, reducing the domain gap for rare but expressive camera angles.
-In addition, we proposed null-pitch conditioning, a data generation step that enhances camera conditioning fidelity when the model is given prompts with conflicting cues.
-We believe our work lays the foundation for physically grounded camera control in video generation, empowering creators to produce dynamic storyboards and convey their visual ideas with greater precision and speed.
+We hope our work lays the foundation for robust camera control under extreme rotations in video generation, empowering creators to produce dynamic storyboards with arbitrary camera orientations and convey their visual ideas with greater precision and speed.
 Acknowledgements.